### v2

Abstract
This paper extends the optimal-trading framework developed [Chriss(2024)] to compute optimal strategies with real-world constraints. The aim of the current paper, as with the previous, is to study trading in the context of multi-player non-cooperative games. While the former paper relies on methods from the calculus of variations and optimal strategies arise as the solution of partial differential equations, the current paper demonstrates that the entire framework may be re-framed as a quadratic programming problem and cast in this light constraints are readily incorporated into the calculation of optimal strategies. An added benefit is that two-trader equilibria may be calculated as the end-points of a dynamic process of traders forming repeated adjustments to each other’s strategy.
Keywords Trading strategies, position-building, game theory, non-cooperative games, equilibrium, equilibrium path, quadratic programming
1 Introduction
Almgren and Chriss developed a robust theory for trading portfolio transactions in Optimal execution of portfolio transactions [^1] See [Almgren and Chriss(2001)] and [Almgren and Chriss(1997)].. That theory and the large body of work that followed is based on the trade-off between transaction costs and risk. The faster one trades, the greater the transaction costs but the more rapidly risk is reduced and vice-versa. A key feature of the framework is the absence of competition from other traders. By contrast [Chriss(2024)] developed a framework for optimal trading that specifically includes competition between two or more traders in the same stock. The nature of the competition studied is non-cooperative, collusion is strictly prohibited.
In the two-trader context, best-response and equilibrium are the central elements of the framework placing trading in a game-theoretic context, specifically non-cooperative games where coordination between traders is strictly prohibited. A best-response strategy is a maximally cost-reducing strategy taking into account an adversary’s strategy. Equilibrium is when two strategies are simultaneously the best-response to one another; [Chriss(2024)] demonstrated that methods from the calculus of variation may be used to derive best-response strategies and two-trader equilibria. The paper additionally contains a definition of $n$-trader symmetric equilibria and explores how empirical aspects of transaction costs and competitive trading influence the structure of optimal strategies.
A limitation of [Chriss(2024)] is that it is limited to unconstrained strategies. In real-world applications it is of considerable interest to find strategies constrained by practical considerations such as position and trading limits, including short-selling. This paper presents computational methods for finding best-response and equilibrium strategies with arbitrary constraints and in the case of two-trader equilibrium sheds additional light on the nature of equilibrium by studying that dynamic path to equilibrium. Therefore if [Chriss(2024)] is about the theoretical foundations of trading as game theory, this paper concerns methods for efficient computation with real-world constraints.
In the immediate following sections we review the definition of trading strategies, best-response strategies and equilibrium. We then show we can approximate the relevant cost functions arbitrarily well by positive-definite quadratic functions and that constraints are linear. With this we are able to translate the minimizations achieved through variational methods to convex quadratic programs in a small number of parameters with linear constraints.
1.1 Definitions and terminology
In this section we define the basic concepts in this paper and introduce standard terminology and notation. Key definitions will be introduced with in separate paragraphs beginning with bold face text.
Position-building: A position-building strategy seeks to acquire a target quantity of stock over a fixed period of time. In this context the motivation for acquiring a position is a catalyst that is expected to cause a revaluation of the stock.
Competition: When two or more traders are simultaneously trading the same stock at the same time they are trading in competition. The competition studied in in [Chriss(2024)] and in this paper is non-cooperative, that is, traders act independently and trading occurs without communication or coordination between traders.
Trading strategies: A trading strategy is a description of how a stock is acquired over time; strategies are represented by twice-differentiable function from the unit interval to the real numbers. The unit interval represents time and the real numbers represent the quantity of stock held at a given time. Trading strategies implicitly trade a specific but unnamed stock. For this paper all trading strategies start at zero and end at the target quantity. Because we will be dealing with two or more traders trading at once, we can always scale one of the traders target quantities to one, which we call a unit strategy. Strategies are denoted by the symbols $a$ or $b$. We use the symbol $\lambda$ for the target quantity of a Non-unit strategy, write $b_{\lambda}$ for the strategy, where $b$ is the associated unit strategy and $b_{\lambda}$ is taken the function $\lambda b$. A feature and limitation of this paper and [Chriss(2024)] is that the strategies considered are static throughout the course of trading. In this way, strategies may be treated as "plays" in games and standard concepts from game theory (see below) may be applied.
Constraints: An unconstrained trading strategy starts and ends at pre-determined quantities but is otherwise free to hold any quantity. at any time. Constrained strategies on the other hand have limits on what quantities of stocks may be held that reflect so-called "real-world" considerations that include, among others, maximum size limits and short-selling constraints. In more detail here are several examples of such constraints:
  • Restrictions on overbuying (see Section 3.2): optimal position-building strategies can significantly relative to the target quantity. In either trader may be subject to restrictions on maximum position size which would lead to a constraint of the form, say, $a(t)\leq 1+\rho$ for all $t$;
  • Channel constraints (see Section 3.3): a trader may wish to confine the trajectory of the trading strategy to a certain channel.
  • End-strategy constraints (see Section 3.4): a trader may wish to have a certain fraction of the target quantity acquired by no later than a certain time.
  • Short-selling constraints (see Section 3.5: quite simply, sometimes a trader building a long position is prohibited from selling short altogether or beyond a certain size.
  • No-sell constraints (see Section 3.6): suppose that $A$ never wants to sell the stock, this translates to a constraint on the first-derivative, namely, $\dot{a}(t)\geq 0$ for all $t$.
Game theory and dynamics: The two key game-theoretical concepts in this paper and [Chriss(2024)] are best-response and equilibrium. A best-response strategy produces the most favorable outcome taking the strategies of other traders as given. This is a completely standard interpretation of the concept of best-response from game theory; similarly equilibrium is when two or more strategies are simultaneously the best-responses to each other [^2] The literature on game theory is extensive and here we mention but a few including the theoretical starting point [Von Neumann and Morgenstern(2007)], Nash’s seminal work introducing the concept that became known as Nash equilibrium in [Nash(1950)], [Nash(1951)], [Nash Jr(1950)] and [Nash(1953)]. See also [Holt and Roth(2004)] for a broader perspective.. As is noted in [Holt and Roth(2004)]:
When the goal is prediction rather than prescription, a Nash equilibrium can also be interpreted as a potential stable point of a dynamic adjustment process in which individuals adjust their behavior to that of the other players in the game, searching for strategy choices that will give them better results.
In the case of position building in competition this is exactly what happens when we introduce computational methods for finding two-player equilibria. In Section 5 we introduce an algorithm for computing two-trader equilibria that in effect replicates this dynamic adjustment process (see 1). The algorithm allows one to understand the entire path leading to equilibrium and helps to explain certain features of equilibria such as in certain cases the cost of trading in equilibrium are higher than non-equilibrium, which can be seen as a direct consequence of the non-cooperative nature of the games.
1.2 Mathematical overview of trading strategies
In this paper we focus primarily on two-trader position building scenarios where there is a unit trading strategy $a$ and a $\lambda$-scaled strategy $b_{\lambda}$, however the most general scenario is when there are $n$-traders trading strategies $a_{i}$, where $i$ represents trader $i$. In both cases strategies satisfy the boundary conditions of vanishing at $t=0$ and acquiring $\lambda_{i}$ units of stock at time $t=1$. In the two-trader setup $a=a_{1}$, $b_{\lambda}=a_{2}$, $\lambda_{1}=1$, $\lambda_{2}=\lambda$.
Market impact: When $n$ traders are in competition it is assumed that their aggregate trading exerts the marginal influence on the price of the stock. A trader seeking to purchase a quantity of stock from time $t=0$ to $t=1$ is concerned with two forms of trading cost. One is the transitory cost associated with paying in excess of the prevailing market price at any given time, and the other is the cumulative cost recorded as the difference between the price prevailing at time $t=0$ and the price paid at the time of a given trade. For strategies $a_{i}$ we model the cost paid by trader $i$ at time $t$ as proportional to:
$\left(\sum_{j}\dot{a}_{j}(t)\right)\dot{a_{i}}(t)+\kappa\left(\sum_{j}a_{j}(t)\right)\dot{a_{i}}(t)$ (1.1)
This cost says that relative to the price at time 00 and executing every trade at the prevailing price at the time of trade depends on the aggregate quantity traded at time $t$ by all traders and the net quantity acquired from time $t=0$ to $t$. The parameter $\kappa$ is called the market impact coefficient and represents the relative contribution between the temporary and permanent impact terms. Many functions in the sequel are parameterized by $\kappa$ and in those cases we will leave $\kappa$ as implicit to de-clutter the notation.
Best-response and equilibrium strategies: [Chriss(2024)] studies finding the strategy $a$ that minimizes its total cost of trading in competition with a strategy $b_{\lambda}$ where the total cost of trading is given by:
$\text{Cost}(a,b_{\lambda})=\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.2)
and therefore the best-response strategy is defined as the solution to:
$\underset{\underset{a(0)=0,a(1)=1}{a:[0,1]\to\mathbf{R}}}{\operatorname{\mathbf{minimize}}}\,\,\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (1.3)
In [Chriss(2024)] a two-trader equilibrium was defined simply as a pair of strategies $(a,b_{\lambda})$ such that $a$ is the best-response to $b_{\lambda}$ and $b_{\lambda}$ is the best-response to $a$.
Best-response and equilibrium strategies. For $a$ trading in competition with a $b_{\lambda}$ the best-response strategy is the one that minimizes total-cost, which can be expressed as the integral of the loss function:
$L_{a}(t;b_{\lambda})=(\dot{a}+\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}$ (1.4)
and we have
$L_{b_{\lambda}}(t;a)=(\dot{a}+\lambda\dot{b}_{\lambda})\lambda\dot{b}_{\lambda}+\kappa(a+\lambda b_{\lambda})\lambda\dot{b}_{\lambda}$ (1.5)
To find the cost-minimizing strategy we use the Euler-Lagrange equation:
$\frac{\partial L_{a}}{\partial a}-\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L_{a}}{\partial\dot{a}}\right)=0$ (1.6)
with the resultant PDE being:
$\ddot{a}=-\frac{\lambda}{2}({\ddot{b}}_{\lambda}+\kappa\dot{b}_{\lambda})$ (1.7)
By the same token, if $\text{Cost}(b_{\lambda},a)=\int_{0}^{1}L_{b_{\lambda}}(a,\kappa)\,\texttt{dt}$ we may also use the Euler-Lagrange equation similarly to derive an expression for the minimal strategy $b_{\lambda}$ trading in competition with $a$:
${\ddot{b}}_{\lambda}=-\frac{1}{2\lambda}(\ddot{a}+\kappa\dot{a})$ (1.8)
Finally in [Chriss(2024)] it was shown that the system of system of equations (1.7) and (1.8) has a pair of solutions $a,b_{\lambda}$ and this means that $a,b_{\lambda}$ are simultaneously the best-responses to one another and is a two-trader equilibrium.
Fourier approximations: We use Fourier series to approximate trading strategies, which will be the computational engine used throughout this paper. Given a trading strategy $a$ we know that the function $a(t)-t$ vanishes at zero and one and therefore the Fourier series for $a-t$ is a sine series. For unit strategies $a,b$ we write $a_{\mathcal{F}},b_{\mathcal{F}}$ for their corresponding Fourier expansions and note the $n$-th coefficient may be computed via integration as follows:
$a_{\mathcal{F}}(t)$ $=t+\sum_{1}^{N}a_{n}\sin(n\pi t),\qquad a_{n}=\int_{0}^{1}(a(t)-t)\sin(n\pi t)\,\texttt{dt}$ (1.9a)
$b_{\lambda\mathcal{F}}(t)$ $=t+\sum_{1}^{N}b_{n}\sin(n\pi t),\qquad b_{n}=\int_{0}^{1}(b(t)-t)\sin(n\pi t)\,\texttt{dt}$ (1.9b)
Observe that $N$ is left implicit and from standard results in Fourier series we know that $a_{\mathcal{F}}\to a$ uniformly as $N\to\infty$ and the coefficients $a_{n}$ decay like $1/n^{2}$ for twice-differentiable functions; identical results hold for $b$. In making these observations we are able to identify the space of unit trading strategies with elements in $\mathbf{R}^{N}$:
$a(t)\sim a_{\mathcal{F}}=t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)\mapsto(a_{1},\dots,a_{N})\in\mathbf{R}^{N}$ (1.10)
This identification will allow us to translate problems about the functions representing trading strategies to minimization problems over sets in $\mathbf{R}^{N}$. We use the next sections to do this.
1.3 Convex combinations of trading strategies and the cost function [^3] This section is new to Version 2. The prior version stated Algorithm 1 without any explanation of why it works.
Consider $0\leq\gamma\leq 1$ and two strategies $b_{\lambda,1}$ and $b_{\lambda,2}$ we can form the convex combination
$b_{\lambda,\gamma}=\gamma b_{\lambda,1}+(1\shortminus\gamma)b_{\lambda,2}$ (1.11)
and then for a strategy $a$ we have the associated cost functions $\text{Cost}(a,\gamma b_{\lambda,1})$, $\text{Cost}(a,(1\shortminus\gamma)b_{\lambda,2})$ and $\text{Cost}(a,b_{\lambda,\gamma})$. It is immediate consequence of the linearity of integration and differentiation and they are related as follows:
$\text{Cost}(a,b_{\lambda,\gamma})=\text{Cost}(a,\gamma b_{\lambda,1})+\text{Cost}(a,(1\shortminus\gamma)b_{\lambda,2})$ (1.12)
The meaning of Eq. 1.12 is as follows. The left-hand side is the cost of $a$ trading in competition with $b_{\lambda,\gamma}$, while the right-hand side is the sum of the cost of $a$ trading in competition with $\gamma b_{\lambda,1}$ and $a$ trading in competition with $(1\shortminus\gamma)b_{\lambda,2}$. Now suppose that $a^{*}$ is optimal versus $b_{\lambda,\gamma}$, that is, that it minimizes $\text{Cost}(a,b_{\lambda,\gamma})$.
2 Fourier series and trading strategies
Best-response strategies are the results of the minimizations in 1.2 involving $\text{Cost}(a,b_{\lambda})$, and similarly for $\text{Cost}(b_{\lambda},a)$. To translate this a minimization problem in $\mathbf{R}^{N}$ we proceed by replacing $a,b$ with $a_{\mathcal{F}},b_{\mathcal{F}}$ in the loss functions $L_{a}$ and $L_{b}$ respectively and then the uniqueness of Fourier series combined with convergence and term-by-term integration results guarantees we can find a suitable approximation for the corresponding cost functions.
2.1 Approximating the cost function with Fourier series
This section provides the computational backbone of this paper. We will show that the cost function is a quadratic function of the Fourier coefficients of the strategies $a,b_{\lambda}$.
Proposition 2.1.
The total cost function $\text{Cost}(a,b_{\lambda})=\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ may be approximated by the a function $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$:
$\text{Cost}_{\mathcal{F}}(a,b_{\lambda})=\frac{1}{2}(2+\kappa)(1+\lambda)+\frac{\pi^{2}}{2}\sum_{n=1}^{N}(a_{n}^{2}+\lambda a_{n}b_{n})n^{2}+\frac{2\kappa\lambda}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{b_{n}-a_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{\lambda b_{m}a_{n}nm}{m^{2}-n^{2}}$ (2.1)
Similarly:
Proposition 2.2 (Approximate cost function for $b$ with respect to $a$).
The approximate cost function $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ is given by the formula:
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\lambda\left\{\frac{1}{2}(2+\kappa)(1+\lambda)\!+\!\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}\!+\!a_{n}b_{n}\right)n^{2}+\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}\!+\!2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}\right\}$ (2.2)
We present the proof of this approximation in Section A.2.
2.2 Properties of the approximate cost function
The approximate cost function $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ in equation (2.1) is quadratic in the Fourier coefficients of $a_{\mathcal{F}}$, and similarly $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ is quadratic in the coefficients of $b_{\mathcal{F}}$. Also, an immediate consequence of Eq. 2.1 and Eq. 2.2 is that the Hessian matrix of both are diagonal and positive definitive, which will imply that the quadratic programs we focus on when adding constraints will be convex. Both cost functions depend implicitly on the number of Fourier coefficients used to approximate $a_{\mathcal{F}}$ and $b_{\mathcal{F}}$ and we can show that as $N\to\infty$ the approximate cost functions converge to the actual cost functions. We state this now and sketch the proof.
Proposition 2.3 (The approximate cost-function converges to the cost-function).
Let $a,a_{\mathcal{F}}$ be as above, noting that $a_{\mathcal{F}}$ depends implicitly on the number of Fourier coefficients $N$. Then
$\text{Cost}_{\mathcal{F},a}(b_{\lambda})\to\text{Cost}(a,b_{\lambda}),\quad N\to\infty$ (2.3)
The analogous statement holds for $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$.
Sketch of proof: To see this we use two key facts about Fourier series. The Fourier coefficients of a twice-differentiable function are unique and the integral of a partial Fourier series is the sum of the integrals of the partial sums and that sum converges to the integral of the function.
2.3 Some useful facts about Fourier series of trading strategies [^5] This section is new to Version 2.
There are a number of basic properties of the Fourier series of trading strategies that we note here.
Convex combinations of trading strategies. Given unit strategies $a_{1},a_{2}$ and $0\leq\gamma\leq 1$, $a_{12}=\gamma a_{1}+(1\shortminus\gamma)a_{2}$ is a convex combination of the strategies. Since $a_{12}(0)=0$ and $a_{12}(1)=1$, it has a Fourier series:
$a_{12,\mathcal{F}}=t+\sum_{n=1}^{\infty}a_{12,n}\sin(n\pi t)$ (2.4)
We can also form the Fourier series of the function $\gamma a_{1}+(1\shortminus\gamma)a_{2}$:
$\gamma a_{1}+(1\shortminus\gamma)a_{2}=t+\sum_{n=1}^{\infty}\left(\gamma a_{1,n}+(1-\gamma)a_{2,n}\right)\sin(n\pi t)$ (2.5)
and it is immediately clear that the series in Eq. 2.5 the expression in Eq. 2.4 and Eq. 2.5 must be equal which immediately implies:
$\gamma a_{1,n}+(1\shortminus\gamma)a_{2,n}=a_{12,n},\qquad\text{for all $n$}$ (2.6)
We will use this fact in Section 4.5.
3 Constraints
In this section we describe some of the types of constraints we can use impose on the strategies we calculate and explain the strategy we employ to make these calculations.
3.1 How to constrain best-response and equilibrium strategies
We outline the steps steps to achieve this as follows. Since each of the two trading strategies $a(t),b(t)$ satisfy the boundary conditions that they vanish at 0 and are 1 at time $t=1$, they may be approximated by sine series in $N$ terms
$a(t)\sim a_{\mathcal{F}}=t+\sum_{n=1}^{N}a_{n}\sin(n\pi t),\qquad b\sim b_{\mathcal{F}}=t+\sum_{n=1}^{N}b_{n}\sin(n\pi t)$ (3.1)
where $a_{n}=2\int_{0}^{1}(a(t)-t)\sin(n\pi t)$, $b_{n}=2\int_{0}^{1}(b(t)-t)\sin(n\pi t)\,\texttt{dt}$, respectively. With this in hand we may identify each strategy with an element of $\mathbf{R}^{N}$:
$a(t)\mapsto(a_{1},\dots,a_{N})\in\mathbf{R}^{N}\qquad b_{\lambda}\mapsto(b_{1},\dots,b_{N})\in\mathbf{R}^{N}$ (3.2)
and therefore we may approximate the loss function $L(t)$ as follows:
$L(t)=(\dot{a}+\dot{b}_{\lambda\mathcal{F}})+\kappa(a+\lambda b_{\lambda})\dot{a}\xmapsto{\hskip 28.45274pt}L_{\mathcal{F}}(t)=(\dot{a}_{\mathcal{F}}+\lambda\dot{b}_{\lambda\mathcal{F}})+\kappa(a_{\mathcal{F}}+\lambda b_{\lambda\mathcal{F}})\dot{a}_{\mathcal{F}}$ (3.3)
By virtue of uniqueness of Fourier series for twice-differentiable functions, $L_{\mathcal{F}}$ is the Fourier series in $N$ terms for $L(t)$. Finally by the fact we can perform term-by-term differentiation of Fourier series we can identify the cost function $\text{Cost}(a,b_{\lambda})$ with an approximation:
[…]
Since the loss function itself is at most degree two, it is clear that $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ is at most a degree-two polynomial in the Fourier coefficients of $a$ and $b$. It is the content of Section 2.1 to derive this function and give an explicit formula for $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ as a quadratic function in the coefficients $a_{1}\dots a_{N}$ and $b_{1}\dots b_{N}$. From this it becomes immediately obvious that $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ is also positive-definite and it is straightforward to show it converges uniformly to $\text{Cost}(a,b_{\lambda})$ as the number of terms $N\to\infty$. It is with this identification that we may translate problems relating to minimizing $\text{Cost}(a,b_{\lambda})$ to quadratic programs $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$.
$\begin{array}[]{lll}\text{(EL)}&\operatorname{\mathbf{minimize}}&\text{Cost}(a,b_{\lambda})\\ &\text{subject to}&\textit{No constraints}\end{array}$ $\begin{array}[]{lll}\text{(QP)}&\operatorname{\mathbf{minimize}}&\text{Cost}_{\mathcal{F}}(a,b_{\lambda})\\ &\text{subject to}&\textit{Constraints}\end{array}$
where the minimizations of $a$ and $b_{\lambda}$ respectively. With the basic approach in-hand we next demonstrate how to how to introduce constraints.
Approximating an interval with a finite set of points: Our aim below is to constrain a function $f$ on an interval $I_{ab}=[a,b]$ well enough that the size of the set of points where a constraint such as $f(t)<g(t)$ for all $t\in I_{ab}$ fails to hold is sufficiently small. Thus we want a set $\mathcal{T}=\{t_{1},\dots,t_{K}\}\subset[a,b]$ such that
$\mu\left(\{t\in I_{ab}\,|\,f(t)>g(t)\,\}\right)<\epsilon$
where $\mu$ is some measure on the real numbers, e.g., Lebesgue measure. We omit the details here, and simply note that for relatively "well-behaved" functions and constraints we can construct sufficient sets for our purposes. In what follows we will simply refer to an approximating set of times with respect to a function (or functions).
Path constraints on trading strategies. Given a function $c:[0,1]\to\mathbf{R}$, a strategy $a$ is said to satisfy the path constraint defined by $c$ if $a(t)\leq c(t)$ for all $t\in[0,1]$. We call $c$ the constraint function. For an approximating set of times $\mathcal{T}=\{t_{1},\dots,t_{K}\}\subset[0,1]$ with respect to $M$ constraint functions $c_{1},\dots,c_{M}$ we consider the following constrained optimization:
$\begin{array}[]{ll}\underset{(a_{1}\cdots a_{n})\in\mathbf{R}^{N}}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(a,b_{\lambda})\\[8.61108pt] \text{subject to}&\quad t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)\leq c_{i}(t),i=1,\dots,M\quad\text{for all $t\in\mathcal{T}$ and $c_{1},\dots c_{M}$}\end{array}$ (3.5)
This is a quadratic program in $N$ variables with $K\times M$ linear constraints. In the rest of this section we demonstrate how to implement a variety of useful constraints using (3.5) but first introduce partial path constraints.
Partial path constraints: We can make a simple modification of path constraints by choosing an approximating set of times $\mathcal{T}^{*}=\{t_{1},\dots,t_{k}\}\subset[t^{*},1]$, where $t^{*}>0$ with respect to some constraint functions $c_{1},c_{2},\dots$ to mean that the constraints only apply to some of the times. We now discuss a selection of path and partial path constraints that arise in practice.
3.2 Overbuying constraint
An constraint is a path constraint $c(t)=1+\rho$ for $\rho>0$ so that $a(t)\leq 1+\rho$ and thus limits the extent to which a best-response strategy can "overbuy" relative to its target quantity of $1$. Figure 1 depicts this in general and Figure 4 shows an example the best-response to a risk-neutral adversary with overbuy constraints.
Figure 1: An example of a over-buying strategy (blue line) exceeds the target quantity of one and therefore violates the overbuy constraint.
3.3 Channel constraints
A trajectory channel constraint seeks to keep the trading trajectory within some bounds defined by a pair of functions $L,U$ such that $L(t)\leq U(t)$. A strategy satisfies the channel constraint if $L(t)\leq a(t)\leq U(t)$ for all $t\in[0,1]$, which is clearly an example of a path constraint. For example, if a traders wants the best-best response to some strategy $b_{\lambda}$ but wants to be no more risk averse than $\sinh(4t)/\sinh(4)$ and take no more risk than a risk-neutral strategy then this is defined by $L(t)=\sinh(4t)/\sinh(4)$ and $U(t)=t$. Figure 2 gives a illustration of this.
Figure 2: The trajectory channel bounded by $\sinh(4t)/\sinh(4)$ below and $t$ above. We may seek a constrained optimization that restricts to strategies that are within this channel and see that the risk averse strategy $y=t^{2}$ is always within the channel while the eager strategy is never within the channel.
3.4 End-strategy constraints
End-strategy constraints are examples of partial path constraints, that is, constraints that only apply to a subset of times, in this case in the interval $[t^{*},1]$ for time $t^{*}>0$. In this case each path constraint pertains to a subset of times $\mathcal{T}^{*}=\{t_{1},\dots,t_{K}\}\subset[t^{*},1]$. In practice, end-strategy constraints require a strategy to exhibit some behavior from time $t^{*}$ to time $t=1$, the start of the "end-strategy". We illustrate the most basic end-strategy constraint, that for some $c<1$, $c\geq a(t)\leq 1$ for $t\in[t^{*},1]$. We depict this constraint in Figure 3, and an illustration is in Figure 3
Figure 3: A visual depiction of a end-strategy constraints with $t^{*}=75\%$ and $c=80\%$. The constraint requires that at time $t=0.75$ the trading strategy is at least 80% completed and does not exceed 100% of the target quantity. The eager strategy (blue line) nor the risk-neutral strategy (red-line) satisfy the constraint. See figure 7 for examples of best-response strategies with end-strategy constraints.
3.5 Short-selling constraints
Short-selling constraints either prohibit short-selling altogether (strict short-selling constraint) or limit it in some way. For both versions of short-selling constraints the problem simple. Choose time points $t_{1},\dots t_{K}$ and $c\leq 0$ and solve:
$\begin{array}[]{ll}\operatorname{\mathbf{minimize}}&\quad\text{Cost}_{\mathcal{F},a}(b_{\lambda})\\[5.38193pt] \text{subject to}&\quad t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)\geq c,\,t\in\mathcal{T}\end{array}$ (3.6)
For strict short-selling constraints set $c=0$ and non-strict constraints $c<0$. This simply places a floor below which holdings cannot go. Note that $c<0$ means since $a(t)$ is trading a unit strategy then $|c|$ expresses the short-selling constraint as a fraction of the target size.
3.6 No-sell constraints
In [Chriss(2024)] it was noted that a strategy shape that arises in the study of position building with competition is the bucket strategy depicted in Figure 1. Suppose that the trader wishes to constrain solutions to optimal best-response strategies to ones that do not trade bucket strategies.
There are two approaches to this. The first is simply to constrain the first derivative of the best response strategy to be non-negative, that is, to insure that $\dot{a}(t)\geq 0$ for all $t\in[0,1]$. We can achieve this using the Fourier approximation of $\dot{a}$:
$\dot{a}_{\mathcal{F}}(t)=1+\sum_{n=1}^{N}a_{n}\,n\pi\cos(n\pi t)$ (3.7)
and for a set of times $t_{1}\cdots t_{K}$ require $\dot{a}_{\mathcal{F}}(t_{k})\geq 0$ for each $k$. This can then be solved using
$\begin{array}[]{ll}\operatorname{\mathbf{minimize}}&\quad\text{Cost}_{\mathcal{F},a}(b_{\lambda})\\[5.38193pt] \text{subject to}&\quad t+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\geq 0,\,t\in\mathcal{T}\end{array}$ (3.8)
No-sell constraints by definition constrain a strategy from selling at any point time along the trajectory, and therefore, will automatically prevent . However, no-sell constraints are much stronger overbuy constraints because they prohibit selling when the strategy is holding less than the target quantity as well as when it holds more. By contrast, overbuy constraints prevent selling as a consequence of the target quantity. If a strategy holds more than the target quantity of $1$ at any time, then selling is required to ultimately satisfy the boundary condition.
4 Numerical examples of strategies with constraints
In the sections that follow we provide numerical examples of best-response and equilibrium strategies with constraints. We run through a collection of strategies that a single adversary might trade. [^7] The code for all of these examples can be found online at https://github.com/vragulin/Optimal_Exec_In_Comp, commit 7bc0c74. To explain this start by noting that every adversary depends on a scaling parameter $\lambda$, those strategies that take into account the market impact coefficient $\kappa$ will depend on $\kappa$ and those arise when, for instance, the strategy is itself a best-response strategy. Finally many strategies are themselves members of parametric families, for example risk-averse strategies depend on a risk aversion parameter, $\sigma$ (see Section 4.3).
4.1 A taxonomy of trading strategies
To organize our thinking we start with a hierarchy of position-building trading strategies, in scenarios where there is a unit strategy $a$ in competition with a $\lambda$-scaled strategy $b_{\lambda}$.
  1. 1.
    Passive: Strategies that are not focused on competition and therefore they are parameterized only by their own parameters and do not require knowledge of the market impact parameter $\kappa$. We include passive strategies in this taxonomy as the foundation of best-response strategies.
    • Risk-averse: These are strategies that prioritize delaying purchases in order to minimize risk at the expense of higher market impact costs. Optimal strategies of this type have the form $b_{\lambda}(t)=\sinh(\sigma t)/\sinh(\sigma)$, where $\sigma>0$ is a risk-aversion parameter.
    • Risk-neutral: These are strategies that balance spreading out trades and the expense of higher risk exposure. The optimal strategy in this category is $b(t)=t$; and
    • Eager: These are strategies that prioritize buying stock quickly in order to trade ahead of an adversary. Strategies in this category have the form $b(t)=\tfrac{e^{-\sigma t}-1}{e^{-\sigma}-1}$ where $\sigma$ is an eagerness parameter.
  2. 2.
    Best-response: are optimal response strategies when trading in competition with some other strategy. The operative question when investigating a best-response strategy is what is it the best response to? These strategies require knowledge of both what the strategy that is being responded to is, and the market impact coefficient $\kappa$. Therefore, as a starting place each passive strategy has a best-response strategy.
    • Risk-averse: Best-response strategies to risk-averse strategies and are discussed in Section 4.3.
    • Risk-neutral: Best-response strategies to risk-neutral strategies and are discussed in Section 4.2.
    • Eager: Best-response strategies to eager strategies were discussed in detail in [Chriss(2024)] discussed again in this paper in 4.4. An noteworthy feature of these is that when the level of eagerness is high, they can respond by selling short; some practitioners find this counter-intuitive and do not want to sell short, so in Section 4.4 we illustrate the no-sell constraint.
  3. 3.
    Equilibrium: These are strategies arise from stable sets of strategies that are pairwise best-responses to one another. In all cases, we focus on a unit strategy that is the equilibrium best-response to one or more other strategies.
    • Two-trader: As described in the introduction an equilibrium is defined by a pair of strategies that are simultaneously the best responses to one another;
    • $n$-trader symmetric: This concerns the situation where $n$ traders all wish to purchase the same quantity of stock.
4.2 Best-response to a risk-neutral adversary
This type of strategy is the best-response to a trader $B$ who trades $\lambda$-scaled ($\lambda\geq 1$) risk-neutral strategy $b_{\lambda}=t$. The corresponding best-response strategy is given by
$a(t)=\left(1+\frac{\lambda\kappa}{4}\right)t-\frac{\lambda\kappa}{4}t^{2}$ (4.1)
Figure 4 shows plots of the best-response to a risk-neutral competitor with some constraints on over-buying.
Figure 4: Best-response strategies $a$ to a risk-neutral adversary trading $b_{\lambda}$ with an overbuying constraint. The plots depict two different overbuying buying constraints (500% and 200% of the target quantity) in dashed red and the unconstrained in solid red.
4.3 Best response to a risk-averse adversary
In this section we illustrate the best-response to a risk-averse trader trading an optimal passive strategy parameterized by $\sigma$. Write $b_{\lambda}(t)$ the $\lambda$-scaled strategy of risk-aversion $\sigma$. In [Chriss(2024)] it was shown that best response to a $\lambda$-scaled risk-averse strategy is given by a strategy $a(t)$ (that also depends on $\lambda$ and $\kappa$, which we leave implicit).
$a(t)=\dfrac{\lambda}{2}(q_{\sigma}(0)-q_{\sigma}(t))+\big{(}1+\dfrac{\lambda}{2}(q_{\sigma}(1)-q_{\sigma}(0))\big{)}\,t$ (4.2)
where $q_{\sigma}(t)$ is an auxiliary function depending on $\sigma$ and $\kappa$:
$q_{\sigma}(t)=\frac{\sinh(\sigma\,t)}{\sinh(\sigma)}+\frac{\kappa}{\sigma}\cdot\frac{\cosh(\sigma\,t)}{\sinh(\sigma)}$ (4.3)
where $q_{\sigma,1}=q_{\sigma}(1)$ and $q_{\sigma,2}=q_{\sigma}(0)$. Figure 5 shows best-response strategies $a(t;{\kappa,\!\lambda,\!\sigma\!})$ from Eq. 4.2.
Figure 5: Best-response strategies $a$ to a risk-averse competitor trading $b_{\lambda}$ with risk-aversion parameter $\sigma$ as in Eq. 4.2 for $\sigma=3$. The results are shown for both unconstrained (solid red) and with the constrained for various values of $\rho$.
4.4 Best response to an eager adversary
Finally we examine the case of the best-response strategy to an eager competitor. For this trader we use the $\lambda$-scaled strategy as follows:
$b_{\lambda}(t)=\frac{e^{-\sigma t}-1}{e^{-\sigma}-1},\qquad\text{eager $\lambda$-scaled, $\sigma$ eagerness}$ (4.4)
as above we solve for the best-response strategy $a(t)$ with boundary conditions $a(0)=0,a(1)=1$ to obtain [^8] In the version of [Chriss(2024)] submitted September 5, 2024 there was an error in the expression for $b_{\lambda}(t;\kappa,\sigma)$, the parameters $\sigma$ and $\lambda$ were taken as a single parameter.:
$b_{\lambda}(t)=\frac{e^{\sigma(-t)}\left(\lambda e^{\sigma}(\sigma-\kappa)-te^{\sigma t}((\lambda+2)\sigma-\kappa\lambda)+e^{\sigma+\sigma t}(-\lambda\sigma+\kappa(\lambda-\lambda t)+(\lambda+2)\sigma t)\right)}{2\left(e^{\sigma}-1\right)\sigma}$ (4.5)
To illustrate best-response to an eager strategy, in Figure 6 we plot $b_{\lambda}(t;\kappa,\sigma)$ for $\sigma=3$ of Eq. 4.5 but as opposed to above we plot with the no-sell constraint (see Section 3.6 due to the high propensity to sell short in these best-responses.
Figure 6: Best-response strategies to an eager competitor $b_{\lambda}(t)$ (in blue) with eagerness parameter $\sigma$=3 as in Eq. 4.5 for all plots. The best-response strategies are in solid and dashed red, where the solid red lines are unconstrained and the dashed red are various constraints on Short-Selling, see Section 3.5. In these plots the trader is not allowed to establish a position below $c$ at any time, i.e. $a(t)\geq c,\,\text{for all }t\in[0,1]$.
Figure 7 shows the best response for the same traders, but with a end-strategy constraint (see Section 3.4).
Figure 7: Best-response strategies to a passive eager competitor trading $b_{\lambda}$ with eagerness $\sigma$ (see Eq. 4.5) for $\sigma=3$ and $\kappa=0.5$ computed with a variety of end-strategy constraints depicted by the grey shaded boxes (see Section 3.4).
4.5 Two-trader equilibrium closed-form solutions
In [Chriss(2024)] two-trader equilibrium strategies were derived as strategies $a$, $b_{\lambda}$ that are simultaneously the best-responses to one another given the market impact coefficient $\kappa$. The results are the equilibrium strategies $a_{\texttt{eq}}$, $b_{\texttt{eq}}$ as follows:
$a_{\texttt{eq}}(t)$ $=-\frac{\left(1-e^{-\tfrac{\kappa t}{3}}\right)\!\!\left(-e^{\tfrac{\kappa}{3}}\left(e^{\tfrac{\kappa}{3}}+e^{2\tfrac{\kappa}{3}}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)}$ (4.6a)
$b_{\texttt{eq}}(t)$ $=\frac{\left(1-e^{-\tfrac{\kappa t}{3}}\right)\!\!\left(e^{\tfrac{\kappa}{3}}\left(e^{\tfrac{\kappa}{3}}+e^{2\kappa/3}+1\right)(\lambda+1)+(\lambda-1)e^{\frac{\kappa t}{3}}+(\lambda-1)e^{\frac{2\kappa t}{3}}+(\lambda-1)e^{\kappa t}\right)}{2\left(e^{\kappa}-1\right)\lambda}$ (4.6b)
If $A$ thinks that $B$ will trade the two-trader equilibrium strategy Eq. 4.6b for a given $\kappa$ and $\lambda$ then $A$’s best-response is to trade $a_{\texttt{eq}}(t)$. If, however, $A$ requires constraints on its strategy then while $A$ can add a constraint and solve for the best-response to $b_{\lambda}$, this will not be an equilibrium strategy, because $B$’s strategy is not the best response to $A$’s constrained best-response to $B$. Nevertheless, Figure 8 depicts these best responses.
Figure 8: Best-response strategies to the two-trader equilibrium strategy $b_{\lambda}$ for various values of $\lambda$; the best-response $a$ is constrained by the channel constraint set at various levels as depicted by the shaded grey areas (see Section 3.3). The channel constraints require that $a$ is always between the risk-neutral lower bound and the specified eager benchmark upper bound (4.4).
5 Computing two-trader equilibrium via iterative quadratic programming
In [Chriss(2024)] it was shown that two-trader equilibrium strategies exist by solving a system with partial differential equations with the Euler-Lagrange equation pertaining to each trader. As a very brief review, the paper derived equations defining best-response strategies for $A$ and $B$ position-building in competition where $B$ trades a $\lambda$-scaled strategy. The loss functions for $A$ and $B$ were as follows:
$L(a,b_{\lambda})$ $=(\dot{a}+\dot{b}_{\lambda\mathcal{F}})\cdot\dot{a}+\kappa(a+\lambda b_{\lambda})\cdot\dot{a}$ (5.1)
$L(b_{\lambda},a)$ $=(\dot{a}+\dot{b}_{\lambda\mathcal{F}})\cdot\lambda\dot{b}_{\lambda}+\kappa(a+\lambda b_{\lambda})\cdot\lambda\dot{b}$ (5.2)
where $L(b_{\lambda},a)$ is the loss function $(\dot{a}+\dot{b}_{\lambda\mathcal{F}})\dot{b}_{\lambda\mathcal{F}}+\kappa(a+\lambda b_{\lambda})\dot{b}_{\lambda\mathcal{F}}$ for $b$ with respect to trading in competition with $a$. Using the Euler-Lagrange equation for $L_{a}$ and $L_{b}$ respectively yields the equilibrium equations:
$\ddot{a}$ $=-\frac{\lambda}{2}(\ddot{b}+\kappa\dot{b})$ (5.3)
$\ddot{b}$ $=-\frac{1}{2\lambda}(\ddot{a}+\kappa\dot{a})$ (5.4)
with the boundary conditions $a(0)=0,b(0)=0,a(1)=1,b(1)=1$. Since Eq. 5.3 and Eq. 5.4 individually describe the conditions under which $a$ and $b$ are respectively paths that are minimal with respect to $L(a,b_{\lambda})$ and $L(b_{\lambda},a)$ respectively, they therefore represent being simultaneously the best response to one another. This is the essence of their being equilibrium strategies. The question is how does one translate this to the quadratic programming setting?
The goal is to find a pair of strategies $(a,b_{\lambda})$ that simultaneously minimizes $\text{Cost}_{\mathcal{F}}(a,b_{\lambda})$ and $\text{Cost}_{\mathcal{F}}(b_{\lambda},a)$ subject to a set of constraints on $a$ and $b_{\lambda}$ respectively. We describe the algorithm now.
Overview of the algorithm. The algorithm may be described as replicating the adjustment dynamic process of traders repeatedly forming best-response strategies to their adversary’s prior strategy. The entire process takes play in $\mathbf{R}^{N}$ where $N>$ is a predetermined number of Fourier coefficients used to approximate each cost function. Choose a number of Fourier coefficients $N$ and an initial guess $b_{\lambda}^{(0)}\in\mathbf{R}^{N}$ representing $b_{\lambda}$. Write $a^{(k)}\in\mathbf{R}^{N}$ and $b_{\lambda}^{(k)}\in\mathbf{R}^{N}$ for respectively the values of Fourier coefficients in $a_{\mathcal{F}}$, $b_{\lambda\mathcal{F}}$ respectively after the $k$-th iteration of an iterative optimization. Write ${\mathfrak{C}}(a;c)$ or ${\mathfrak{C}}(b;d)$ for path or partial path constraints for strategy $a$ or $b_{\lambda}$. Write $\text{Cost}_{\mathcal{F}}(a;b_{\lambda}^{(k)})$ to mean the cost function for $a$ with the value for the strategy $b$ fixed as the $k$-th guess, and similarly for $\text{Cost}_{\mathcal{F}}(b;a^{(k)})$. Then we have two separate minimization problems:
$\begin{array}[]{cl}\underset{(a_{1}\cdots a_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(a;c)\end{array}$ (5.5) $\begin{array}[]{cl}\underset{(b_{1}\cdots b_{N})}{\operatorname{\mathbf{minimize}}}&\quad\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(k)})\\[8.61108pt] \text{subject to}&\quad{\mathfrak{C}}(b;d)\end{array}$ (5.6)
The algorithm works by solving a pair of problems in each iteration. To start, in Step 1(i) we solve Eq. 5.5 for $\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(0)})$ which yields an $\operatorname{\mathbf{arg-min}}$ of $a^{(1)}$. Then in Step 1(ii) we solve (5.6) for $\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(1)})$ to yield an $\operatorname{\mathbf{arg-min}}$ of $b^{(1)}$:
Step 1(i):Minimize Cost(a,bλ(0)) s.t. (a;c)Produces: a(1)Step 1(ii):Minimize Cost(bλ,a(1)) s.t. (b;d) Produces: bλ(1) $\begin{array}[]{lll}\textbf{Step 1(i):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(0)})$ s.t. ${\mathfrak{C}}(a;c)$}&\quad\textbf{{Produces}: $a^{(1)}$}\\ \textbf{Step 1(ii):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(1)}$) s.t. ${\mathfrak{C}}(b;d)$ }&\quad\textbf{Produces: $b_{\lambda}^{(1)}$ }\end{array}$ (5.7)
Then the for $k>1$, the $k$-th step repeats:
Step k(i):Minimize Cost(a,bλ(k1)) s.t. (a;c)Produces: a(k)Step k(ii):Minimize Cost(bλ,a(k)) s.t. (b;d) Produces: bλ(k) $\begin{array}[]{lll}\textbf{Step k(i):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(a,b_{\lambda}^{(k\shortminus 1)})$ s.t. ${\mathfrak{C}}(a;c)$}&\quad\textbf{{Produces}: $a^{(k)}$}\\ \textbf{Step k(ii):}&\quad\text{Minimize $\text{Cost}_{\mathcal{F}}(b_{\lambda},a^{(k)}$) s.t. ${\mathfrak{C}}(b;d)$ }&\quad\textbf{Produces: $b_{\lambda}^{(k)}$ }\end{array}$ (5.8)
The result of the optimization is a pair of points $(a^{*},b^{*})\in\mathbf{R}^{N}\times\mathbf{R}^{N}$ that correspond to the reconstructed functions which we identify by the same symbols:
$\begin{split}a^{*}&=(a_{1},\dots,a_{N})\mapsto a^{*}(t)=t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)\\ b^{*}&=(b_{1},\dots,b_{N})\mapsto b^{*}(t)=t+\sum_{n=1}^{N}b_{n}\sin(n\pi t)\end{split}$ (5.9)
In practice we introduce a parameter $0<\gamma\leq 1$ that allows one to step "part of the way" toward the new function at each step. We show the complete algorithm in 1 below.
Algorithm 1 (Alternating two-trader equilibrium)
1:  initialize: $b^{(0)},{\mathfrak{C}}(a;c),{\mathfrak{C}}(b_{\lambda};d)$, $k=1$, $N$, $10<\gamma<$ 1.
2:  while not converged do
3:     Step k(i):
4:      Compute: $a=\operatorname{\mathbf{arg-min}}_{(a_{1}\cdots a_{N})}\text{Cost}_{\mathcal{F}}\big{(}a,b_{\lambda}^{(k\shortminus 1)}\big{)}\operatorname{\mathrm{s.t.}}\,{\mathfrak{C}}(a;c)$, see (5.5)
5:     $\quad a^{(k)}\leftarrow\gamma a+(1\shortminus\gamma)a^{(k\shortminus 1)}$
6:     Step k(ii):
7:      Compute: $b=\operatorname{\mathbf{arg-min}}_{(b_{1}\cdots b_{N})}\text{Cost}_{\mathcal{F}}\big{(}b_{\lambda},a^{(k)}\big{)}\,\operatorname{\mathrm{s.t.}}\,{\mathfrak{C}}(b_{\lambda};d)$, see (5.6);
8:      $b_{\lambda}^{(k)}\leftarrow\gamma b+(1-\gamma)b_{\lambda}^{(k\shortminus 1)}$
9:     Update: $k\leftarrow k+1$
10:  end while
11:  output: $a^{*}=(a_{1},\dots,a_{N}),b^{*}=(b_{1},\dots,b_{N})$; $a^{*}(t)=t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)$ and $b^{*}=t+\sum_{n=1}^{N}b_{n}\sin(n\pi t)$.
Note that in each successive step in 1 optimize by minimizing the cost function trading not versus the prior best-response strategy but rather versus a convex combination of the prior best-response and current best-response’s Fourier coefficients. As we saw in Section 2.3, the convex combination of the strategy’s Fourier coefficients is an approximation of the same convex combination of strategies. We find that in practice this leads to significantly better convergence. In what follows we overload the term best-response to include these interpolations. See Figure 15 for an illustration of the impact of $\gamma$ on the equilibrium path.
Intuition. Algorithm 1 has an intuitive interruption which explains why it will always converge to the two-trader equilibrium. After initialization, each step is simply the best-response to the other trader’s best-response strategy. For example, to compute $a^{(k)}$ is to find the best-response to $b_{\lambda}^{(k\shortminus 1)}$. The algorithm proceeds by successive best-responses to the last trader’s best-response strategy until there is no more advantage to be gained, precisely the definition of equilibrium. Because this is a non-cooperative game, each trader can only react to the current best-response strategy as there is no available information on what either trader intends to do next. In Section 6.1 we demonstrate that this can lead to cases where the cost of equilibrium are actually higher than non-equilibrium. We do this through the use of state-space diagrams and we use this to provide additional insight into the nature of the equilibria we observe.
6 Numerical examples of two-trader equilibrium
In this section we provide numerical examples to illustrate using Algorithm 1 to calculate two-trader equilibria. As usual there is a unit trader with strategy $a$ and a $\lambda$-scaled trader with strategy $b_{\lambda}$. We look at scenarios where $\kappa$ ranges from $1$ to $25$ with various values for $\lambda$. In the first part below we look at unconstrained two-trader equilibria. In this part we compare the output of Algorithm 1 to the results given by the closed-form solutions in (4.6a) and (4.6b).
6.1 Computing equilibria and the dynamic path to equilibrium
We begin by studying two-trader equilibria with no constraints to achieve two purposes. With no constraints we can compare the optimization results to the closed form solutions in (4.6a) and (4.6b); and we can study the path to convergence build up from each step in Algorithm 1. This provides insight into the nature of the dynamic process of adjustments that leads to convergence.
To provide a measure of the quality of the results, we use the $L_{2}$ norm of the difference between the Fourier approximation and the closed form solution. For $\kappa$ and $\lambda$ let $(a_{1},\dots,a_{N})$ be the Fourier coefficients of the result of Algorithm 1 for strategy $a$. Then let $a^{*}(t)=t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)$ be the corresponding reconstructed function; and let $a_{\texttt{eq}}(t)$ be the two-trader equilibrium strategy for the same $\kappa,\lambda$ in Eq. 4.6a. Write $\|a^{*}-a_{\texttt{eq}}\|_{2}$ for the $L_{2}$ norm of the difference between $a^{*}$ and $a_{\texttt{eq}}$:
$\|a^{*}-a_{\texttt{eq}}\|_{2}=\left(\int_{0}^{1}\left(a^{*}(t)-a_{\texttt{eq}}(t)\right)^{2}\,\texttt{dt}\right)^{1/2}$ (6.1)
We will use this as our measure of the quality of the solution $a^{*}$ versus the closed-form solution $a_{\texttt{eq}}$. Similarly we can form $b^{*}$ and use $\|b^{*}-b_{\texttt{eq}}\|_{2}$ to measure the quality of the solution of $b^{*}$.
Plots: Figures 9 to 11 provide three illustrations of solving the unconstrained two-trader equilibrium using Algorithm 1. Each optimization starts with the initial guess $a^{(0)}=t$, the risk-neutral strategy and the first iteration is $b_{\lambda}^{(1)}$ the best-response to this strategy. Unless otherwise stated, these plots $\gamma=0.8$ (see 1). We use a lower value when convergence is an issue. The upper left-hand plot of each shows the trading strategies for $a$ and $b_{\lambda}$ on top of one another; the upper right-hand plot shows the difference between the closed-form and approximate solutions for both the initial iteration and the final iteration of the optimization. For example, in Figure 9 ($\lambda=5$ and $\kappa=1$) the right hand plot shows in solid red the plot of $b^{(1)}(t)-b_{\texttt{eq}}(t)$ which shows that at approximately $t=0.5$ the equilibrium strategy strategy holds approximate 0.35 units more than the initial guess (the risk-neutral strategy). The initial guess for $a$ is a lot closer in absolute terms (dark blue) but with $\lambda=5$ this scale difference makes sense.
Equilibrium paths and consequences of no-collusion: As noted earlier Algorithm 1 represents the dynamic path to the equilibrium strategy, showing successive best-responses to the prior best-response strategy. In cases where two traders are in competition and trading similar quantities of stock the results reveal an important consequence of prohibitions on collusion. The conclusion is that in certain instances where two traders are trading approximately the same quantity, the costs to each is lower when they both trade the risk-neutral than when they trade the equilibrium strategy. However, because the traders cannot coordinate their trading they each must successively fashion the best-response to the other’s best response and inevitably pay more. This is impossible to avoid without some form of collusion.
We can see this in Figure 11 which displays data for the unconstrained two-trader equilibrium for the case where $\lambda=1$ and $\kappa=5$ and specifically $\gamma=0.2$. The lower left-hand plot shows the state-space diagram for the optimization. The axes of the plot are the cost of trading for $a$ and $b$ respectively and it plots $(a^{(k)},b^{(k\shortminus 1)})$, the result of step k(i), and $(a^{(k)},b^{(k)})$, the result of step k(ii). The point labeled init guess shows the costs of the pre-initialization strategies, that is, when both $a$ and $b$ trade the risk-neutral strategy. Both costs are approximately 27. The initial guess is connected by a line to a second point whose coordinates are approximately $(22,32)$. This point represents the pair of strategies consisting of $a$, which is a convex combination of the best-response to $b$ trading the risk-neutral strategy and the risk-neutral strategy (see the use of $\gamma$ in Algorithm 1). Clearly this step decreases $a$’s total cost of trading while it increases $b$’s. However, the next iteration, which is $b$ trading the best response to $a$ from the prior iteration, decreases the total cost of trading for $b$ while increasing the cost for $a$. And it exactly as such. The net result is that both traders end up paying more than simply trading the risk-neutral strategies. An obvious conclusion is that if the two-traders were allowed to collude they would be able to conclude that both trading the risk-neutral strategy would be mutually beneficial but as with many two-person non-cooperative games, such as prisoners dilemma. A similar phenomena is also show in the right two plots in Figure 12.
6.2 Plots of unconstrained two-trader equilibrium
In the following plots we examine two-trader equilibria for various values of $\kappa$ and $\lambda$ with $\gamma=0.2$ in all cases. At this stage we have no investigated what the optimal value for $\gamma$ is but illustrate in the sequel that the algorithm proposed has instances that converge with $\gamma<1$ but not with $\gamma=1$.
Figure 9: Two-trader equilibria strategies $a^{*},b_{\lambda}^{*}$ compared with the closed-form solutions $a_{\texttt{eq}},b_{\texttt{eq}}$ in (4.6a) and (4.6b) for $\kappa=1$ and $\lambda=5$ using $N=20$ Fourier terms. The upper left-hand plot shows $a^{*}$ traded along with $a_{\texttt{eq}}$ and $b^{*}$ along with $b_{\texttt{eq}}$, as well as the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The upper right-hand shows the difference between the reconstructed functions and theoretical equilibrium functions at the first and last iterations of the optimization. The state-space diagram in the lower-left plot shows the total cost of trading for $a$ and $b_{\lambda}$ directly prior to Step 1(i) when both traders trade a risk-neutral strategy. The plot shows that the solution (the equilibrium strategy) benefits $a$, lowering the total cost trading from approximately 9 to 8.2, while increasing $b_{\lambda}$’s modestly 45 to approximately 46.2. The solver converged in 12 iterations and $\gamma=0.8$.
Figure 10: Two-trader equilibria solutions compared with the closed-form for $\lambda=20,\kappa=6$ and $N=25$ Fourier series terms for Algorithm 1 with $\gamma=0.8$ and Algorithm 1 converged in 21 iterations. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations.. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. In this case the $b_{\lambda}$ trades 20 times as much as $a$ and $a$’s initial best-response to $b_{\lambda}=\lambda\cdot t$ already substantially reduces its cost while increasing $b_{\lambda}$’s.
Figure 11: Unconstrained two-trader equilibria solutions compared with the closed-form for $\lambda=1,\kappa=25$ and $N=35$ terms in the Fourier series approximations for Algorithm 1 with $\gamma=0.2$ and the algorithm converged in 62 iterations. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. Of interest is that the total cost of trading is higher for both traders at equilibrium than it would be if they both traded a risk-neutral strategy as they do at the point labeled init guess. That is, they would both be better off if they could coordinate. It’s also interesting that the convergence is much slower for higher $\gamma$, for example if we use the same value of $\gamma=0.8$ as in the previous example, the solver does converge within tolerance even after 100 iterations.
Ensemble plots: In Figures 13 to 17 we depict "ensembles" of plots of two-trader equilibrium to demonstrate how a single parameter’s change influences the optimization results. For example, as explained in Algorithm 1 it can be helpful to advance in each step using convex combinations of prior- and next-step strategies. We have not explored how to optimally set $\gamma$ but in Figure 16 we demonstrate for the case $\kappa=25$ and $\lambda=1$ the algorithm diverges when $\gamma=1$ and converges when $\gamma\leq 0.6$.
Figure 12: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with the values of $\gamma$ selected to ensure quick convergence (i.e. $\gamma=0.8$ for the left two plots, and $\gamma=0.2$ for the right three). The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The number of terms $N$ for each set of $\lambda$ and $\kappa$ is chosen to ensure a given level of approximation accuracy, which we measure by the $L_{2}$ norm of the difference between the exact analytic and the approximated trading strategy (shown above each top chart). The state-space diagram in the bottom row plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labelled init guess), which represents both traders trading the risk-neutral strategy. Of interest is for the two rightmost charts the total cost of trading is higher for both traders than it would be if they both traded a risk-neutral strategy as in initiation – they would both be better off, but because collusion is prohibited, they inevitably both end up trading the equilibrium strategies, which are more expensive than trading the risk-neutral strategies.
Figure 13: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with the values of $\gamma$ selected to ensure quick convergence (i.e. $\gamma=0.8$ for the left two plots, and $\gamma=0.2$ for the right three). The sets of parameters $\lambda$, $\kappa$, $\gamma$ is the same as in the previous chart 12, but the number of Fourier terms is fixed at lower $N=10$ for all charts. Both the $L_{2}$ norms of the approximation error and visual inspection illustrate the dependence of the approximation accuracy on the number of Fourier terms $N$. For example, for $\kappa=20$, $\lambda=1$ parameters (second chart from the right) the $L_{2}$ approximation error for $N=10$ is almost 12 times higher than for $N=30$ (0.0025 vs. 0.0002).
Figure 14: This plot demonstrates the impact of a the number of Fourier coefficients $N$ used in forming the approximate cost functions (see Propositions 2.1 and 2.2) for the case $\kappa=20,\lambda=5$ and $\gamma=0.5$.
Figure 15: Unconstrained two-trader equilibrium for $\kappa=5$, $\lambda=1$ for $\gamma$ increasing from left to right from 0.2 to 1.0; higher $\gamma$ decreases the number of iterations required for convergence, while higher $\gamma$ eventually causes the solver to diverge.
Figure 16: A case where Algorithm 1 diverges when $\gamma$ is too large.
6.3 Two-trader equilibria with constraints
As a final set of numerical demonstrations we calculate two-trader equilibria with constraints using Algorithm 1 to calculate two-trader equilibria with various constraints.
Figure 17: Two-trader constrained equilibria solutions compared with the unconstrained equilibrium closed-form for a cross-section of parameters using Algorithm 1 with $\gamma=0.2$. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}_{\lambda}$ in Eq. 5.9, representing constrained equilibrium trading strategies, subject to constraints on overbuying and short selling, see 3.2 and 3.5. For reference the chart also includes the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The number of terms $N$ for each set of $\lambda$ and $\kappa$ is chosen to ensure high approximation accuracy. The state-space diagram in the bottom row plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labelled init guess), which represents both traders trading the risk-neutral strategy. As the relative size of trader B, $\lambda$, increases, trader A tries to build up progressively greater positions to benefit from the permanent market impact, $\kappa$, until the overbuy constraint is hit. Of interest is that for the left column of charts the total cost of trading is higher for both traders than it would be if they both traded a risk-neutral strategy as in initiation – they would both be better off, but because collusion is prohibited, they inevitably both end up trading the constrained equilibrium strategies, which are more expensive than trading the risk-neutral strategies. In all other columns, the smaller trader A reduces her trading costs at the expense of trader B, whose costs increase.
Appendix A Proofs of cost function approximations
For completeness this section proves the expressions for the cost function approximations given in Propositions 2.1 and 2.2. These proofs are largely a consequence of the trigonometric identities in Section A.3 and basic calculus.
A.1 Proof of Proposition 2.1
To prove this we proceed by noting the total cost function may be expressed as the sum of four integrals by expanding the terms of the loss function $(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}$:
$\text{Cost}(a,b_{\lambda})$ $=\int_{0}^{1}(\dot{a}+\lambda\dot{b}_{\lambda})\dot{a}+\kappa(a+\lambda b_{\lambda})\dot{a}\,\texttt{dt}$ (A.1)
$=\int_{0}^{1}\dot{a}\dot{a}\,\texttt{dt}+\int_{0}^{1}\lambda\,\dot{b}_{\lambda}\dot{a}\,\texttt{dt}+\int_{0}^{1}\kappa\,a\dot{a}\,\texttt{dt}+\int_{0}^{1}\kappa\lambda\,b_{\lambda}\dot{a}\,\texttt{dt}$ (A.2)
$=\int_{I}+\int_{II}+\int_{III}+\int_{IV}$ (A.3)
where each term is determined by the appropriate Fourier expansions:
The first term $\int_{I}=\int_{0}^{1}\dot{a}\dot{a}\,\texttt{dt}$: We substitute $\dot{a}_{\mathcal{F}}$ for $\dot{a}$
$\int_{I}$ $=\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{a}_{F}\,\texttt{dt}$ (A.4)
$=\int_{0}^{1}\left(1+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\right)\left(1+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\right)\,\texttt{dt}$ (A.5)
$=\int_{0}^{1}1+2\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)+\Big{(}\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\Big{)}^{\!2}\,\texttt{dt}$ (A.6)
$=\int_{0}^{1}1\,\texttt{dt}+2\sum_{n=1}^{N}a_{n}\int_{0}^{1}n\pi\cos(n\pi t)\,\texttt{dt}+\sum_{n=1}^{N}\sum_{m=1}^{N}\int_{0}^{1}a_{n}a_{m}n^{2}m^{2}\pi^{2}\cos(n\pi t)\cos(m\pi t)\,\texttt{dt}$ (A.7)
$=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}$ (A.8)
The final equality follows from the identities (1.9b) and the trigonometric integral identity (A.53c), respectively. Summarizing:
$\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{a}_{F}\,\texttt{dt}=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}$ (A.9)
The second term $\int_{II}$ is given by $\int_{0}^{1}\lambda\,\dot{b}\dot{a}\,\texttt{dt}$ and substituting $\dot{a}_{\mathcal{F}},\dot{b}_{\lambda\mathcal{F}}$ for $a$ and $b$ respectively:
$\int_{II}$ $=\int_{0}^{1}\lambda\,\dot{b}_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}$ (A.10)
$=\lambda\int_{0}^{1}\Big{(}1+\sum_{n=1}^{N}b_{n}n\pi\cos(n\pi t)\Big{)}\Big{(}1+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\Big{)}\,\texttt{dt}$ (A.11)
$=\lambda\int_{0}^{1}1+\sum_{n=1}^{N}b_{n}n\pi\cos(n\pi t)+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)+\sum_{n=1}^{N}b_{n}n\pi\cos(n\pi t)\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\,\texttt{dt}$ (A.12)
$=\int_{0}^{1}1\,\texttt{dt}+2\sum_{n=1}^{N}(b_{n}+a_{n})\int_{0}^{1}n\pi\cos(n\pi t)\,\texttt{dt}+\sum_{n=1}^{N}\sum_{m=1}^{N}\int_{0}^{1}b_{n}a_{m}n^{2}m^{2}\pi^{2}\cos(n\pi t)\cos(m\pi t)\,\texttt{dt}$ (A.13)
$=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}$ (A.14)
Where the final equality follows from the identities (A.53b) and (A.53d) of Section A.3, respectively. Summarizing:
$\int_{0}^{1}\lambda\,\dot{b}_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}$ (A.15)
[…]
$\int_{III}$ $=\int_{0}^{1}\kappa\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}\,\texttt{dt}$ (A.16a)
$=\int_{0}^{1}\kappa\,\Big{(}1+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\Big{)}\Big{(}t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)\Big{)}\,\texttt{dt}$ (A.16b)
$=\kappa\int_{0}^{1}t+\sum_{n=1}^{N}a_{n}\sin(n\pi t)+t\cdot\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\sum_{n=1}^{N}a_{n}\sin(n\pi t)\,\texttt{dt}$ (A.16c)
$=\kappa\!\!\left(\int_{0}^{1}t\,\texttt{dt}+\!\!\sum_{n=1}^{N}\int_{0}^{1}a_{n}\sin(n\pi t)\,\texttt{dt}+\!\!\sum_{n=1}^{N}\int_{0}^{1}a_{n}n\pi\cos(n\pi t)\,\texttt{dt}+\!\!\int_{0}^{1}\sum_{m,n=1}^{N}a_{n}a_{m}n\pi\cos(n\pi t)\sin(m\pi t)\,\texttt{dt}\right)$ (A.16d)
To integrate the last term Eq. A.16d we use the identity Eq. A.53e from Section A.3:
$\kappa\int_{0}^{1}\sum_{m,n=1}^{N}a_{n}a_{m}n\pi\cos(n\pi t)\sin(m\pi t)\,\texttt{dt}$ $=\kappa\!\!\sum_{n,m=1}^{N}\int_{0}^{1}a_{n}a_{m}n\pi\cos(n\pi t)\sin(m\pi t)\,\texttt{dt}$ (A.17)
$=\kappa\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\pi\frac{2m}{\pi(m^{2}-n^{2})}$ (A.18)
$=2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$ (A.19)
and we note that this term is zero because for every pair $(m,n),n+m\text{odd}$, the is a corresponding pair $(m,n)$ which cancels. This leads to Eq. A.16d becoming
$\frac{\kappa}{2}-\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}nm\cdot\frac{a_{n}a_{m}}{m^{2}-n^{2}}$ $=\frac{\kappa}{2}$ (A.20)
Summarizing we have the following expression for $\int_{III}$:
$\int_{0}^{1}\kappa\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}=\frac{\kappa}{2}$ (A.21)
The fourth term $\int_{IV}$ corresponds to $\int_{0}^{1}\kappa\lambda\,b\dot{a}\,\texttt{dt}$ and we proceed by substituting $b_{\lambda\mathcal{F}}$ for $b$ and $a_{\mathcal{F}}$ for $a$:
$\int_{IV}$ $=\int_{0}^{1}\kappa\lambda b_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}$ (A.22)
$=\kappa\lambda\int_{0}^{1}\Big{(}t+\sum_{n=1}^{N}b_{n}\sin(n\pi t)\Big{)}\Big{(}1+\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\Big{)}\,\texttt{dt}$ (A.23)
$=\kappa\lambda\int_{0}^{1}t+t\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)+\sum_{n=1}^{N}b_{n}\sin(n\pi t)+\sum_{n=1}^{N}b_{n}\sin(n\pi t)\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\,\texttt{dt}$ (A.24)
To integrate the last term we proceed analogously to what we did with $\int_{III}$:
$\kappa\lambda\int_{0}^{1}\sum_{n=1}^{N}b_{n}\sin(n\pi t)\sum_{n=1}^{N}a_{n}n\pi\cos(n\pi t)\,\texttt{dt}$ $=\kappa\lambda\sum_{n,m=1}^{N}\int_{0}^{1}a_{n}b_{m}n\pi\cos(n\pi t)\sin(m\pi t)\,\texttt{dt}$ (A.25)
$=2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\frac{m}{m^{2}-n^{2}}$ (A.26)
We can see that the second and third terms of this integral do not cancel and instead become sums of terms $\frac{2(b_{n}-a_{n})}{n\pi}$ and so the final expression for $\int_{IV}$ is as follows:
$\int_{0}^{1}\kappa\lambda b_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}=\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{a_{m}b_{n}}{m^{2}-n^{2}}\right)$ (A.27)
Computing the cost function as the sum $\int_{I}+\int_{II}+\int_{III}+\int_{IV}$: Finally we arrive at the cost function via summing the previous results. Summarizing we have:
$\int_{I}$ $=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}$ (A.28)
$\int_{II}$ $=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}$ (A.29)
$\int_{III}$ $=\frac{\kappa}{2}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$ (A.30)
$\int_{IV}$ $=\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)$ (A.31)
Collecting the constant terms we have:
$1+\lambda+\frac{\kappa}{2}+\frac{\kappa\lambda}{2}$ $=1+\lambda+\frac{\kappa(1+\lambda b_{\lambda})}{2}$ (A.32)
$=\frac{(2+\kappa)(1+\lambda b_{\lambda})}{2}$ (A.33)
Next combining the non-constant terms of integrals I and II and the first sum of integral IV we obtain:
$\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}(a_{n}^{2}+\lambda a_{n}b_{n})+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}$ (A.34)
and finally combining the summation in integral III and the last summation integral IV we have:
$2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}$ (A.35)
Putting the three terms together we arrive at:
$\text{Cost}_{\mathcal{F},a}(b_{\lambda})=\frac{(2+\kappa)(1+\lambda)}{2}+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}(a_{n}^{2}+\lambda a_{n}b_{n})+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{\lambda b_{m}a_{n}\,n\,m}{m^{2}-n^{2}}$ (A.36)
Rearranging the terms we obtain the proof of Proposition 2.1.
A.2 Proof of Proposition 2.2
The derivation of the formula proceeds almost identically with the prior one as follows:
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ $=\int_{0}^{1}(\dot{a}_{\mathcal{F}}+\lambda\dot{b}_{\lambda\mathcal{F}})\lambda\dot{b}_{\lambda\mathcal{F}}+\kappa(a_{\mathcal{F}}+\lambda b_{\lambda\mathcal{F}})\lambda\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}$ (A.37)
$=\lambda\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}+\lambda^{2}\int_{0}^{1}\dot{b}_{\lambda\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}+\kappa\lambda\int_{0}^{1}a_{\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}+\kappa\lambda^{2}\int_{0}^{1}b_{\lambda\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}$ (A.38)
$=\int_{I}^{ba}+\int_{II}^{ba}+\int_{III}^{ba}+\int_{IV}^{ba}$ (A.39)
Recall:
$\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{a}_{F}\,\texttt{dt}$ $=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}$ (A.40a)
$\lambda\int_{0}^{1}\,\dot{b}_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}$ $=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}$ (A.40b)
$\kappa\int_{0}^{1}\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}\,\texttt{dt}$ $=\frac{\kappa}{2}+2\kappa\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$ (A.40c)
$\kappa\lambda\int_{0}^{1}b_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}$ $=\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)$ (A.40d)
With the above identities we can provide formulas for $\int_{I}^{ba},\dots,\int_{IV}^{ba}$.
Computing integral I: To start note that $\int_{I}^{ba}$ has exactly the same form as $\int_{II}$, Eq. A.40b except the roles of $\dot{a}_{\mathcal{F}}$ and $\dot{b}_{\lambda\mathcal{F}}$ reversed. Since the latter is symmetric in $a_{\mathcal{F}}$ and $b_{\lambda\mathcal{F}}$. The form of $\int_{II}$ is
$\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{a}_{F}\,\texttt{dt}=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}\$
so
$\int_{I}^{ba}=\lambda\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}$ (A.41)
Computing integral II: Next $\int_{II}^{ba}$ involves $\dot{b}_{\lambda\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}$ which is the same as $\int_{I}$, Eq. A.40a except for switching $b_{\lambda\mathcal{F}}$ for $a_{\mathcal{F}}$ and a factor of $\lambda^{2}$. The form of $\int_{I}$ is given by:
$\int_{0}^{1}\dot{a}_{\mathcal{F}}\dot{a}_{F}\,\texttt{dt}=1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}^{2}$
so
$\int_{II}^{ba}=\lambda^{2}\int_{0}^{1}\dot{b}_{\lambda\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}+=\lambda^{2}\left(1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}b_{n}^{2}\right)$ (A.42)
Computing integral III: Next $\int_{III}^{ba}$ involves $a_{\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}$ which is the same as $\int_{IV}$, Eq. A.40d, except for switching the roles of $b_{\lambda\mathcal{F}}$ for $a_{\mathcal{F}}$. The form of $\int_{IV}$ is:
$\kappa\lambda\int_{0}^{1}b_{\lambda\mathcal{F}}\dot{a}_{\mathcal{F}}\,\texttt{dt}=\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)$
And therefore:
$\int_{III}^{ba}=\kappa\lambda\int_{0}^{1}a_{\mathcal{F}}\dot{b}_{\lambda\mathcal{F}}\,\texttt{dt}+=\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+2\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)$ (A.43)
Computing integral IV: Finally $\int_{IV}$ involves $b_{\lambda\mathcal{F}}$ and $\dot{b}_{\lambda\mathcal{F}}$ which mirrors Eq. A.40c with $b_{\lambda\mathcal{F}}$ replacing $a_{\mathcal{F}}$ and adding a factor of $\lambda^{2}$. Therefore since $\int_{IV}$ has the form:
$\kappa\int_{0}^{1}\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}\,\texttt{dt}=\frac{\kappa}{2}+2\kappa\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$
we have:
IVba=κλ201bλb˙λdt
=κλ22+2κλ2n,m=1n+moddNbnbmnmm2n2
[…]
Putting it all together: To compute $\text{Cost}_{\mathcal{F},b_{\lambda}}(a)$ we compute the sum $\int_{I}^{ba}+\int_{II}^{ba}+\int_{III}^{ba}+\int_{IV}^{ba}$:
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}+\\ &\quad\lambda^{2}\left(1+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}b_{n}^{2}\right)+\\ &\quad\kappa\lambda\left(\frac{1}{2}+\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+2\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)+\\ &\quad\frac{\kappa\lambda^{2}}{2}+2\kappa\lambda^{2}\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\end{split}$ (A.45)
Collecting constants:
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\\ &\quad\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}a_{n}b_{n}+\lambda^{2}\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}b_{n}^{2}+\\ &\quad\kappa\lambda\left(\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+2\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}\right)+\\ &\quad 2\kappa\lambda^{2}\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\end{split}$ (A.46)
Organizing:
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+\\ &\quad 2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+2\kappa\lambda^{2}\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\end{split}$ (A.47)
Simplifying
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n\pi}+\\ &\quad 2\kappa\lambda\left(\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+\lambda\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\right)\end{split}$ (A.48)
finally noting that the last is zero due to cancellations, we have: the last expression in Eq. A.48 becomes
$2\kappa\lambda\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}$ (A.49)
Next we note:
$\begin{split}\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}&=\lambda\left(1+\lambda++\frac{\kappa}{2}+\frac{\kappa\lambda}{2}\right)\\ &=\lambda\frac{2(1+\lambda)+\kappa(1+\lambda)}{2}\\ &=\frac{\lambda(1+\lambda)(2+\kappa)}{2}\end{split}$ (A.50)
so then
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\frac{\lambda(1+\lambda)(2+\kappa)}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{(a_{n}-b_{n})}{n\pi}+2\kappa\lambda\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}$ (A.51)
The conclusion after some re-arrangement and noting that the last term with $\lambda b_{n}b_{m}nm$ cancel so that we arrive at the formula conform to $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ in Proposition 2.2:
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\lambda\left\{\frac{1}{2}(2+\kappa)(1+\lambda)\!+\!\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}+a_{n}b_{n}\right)n^{2}\!+\!\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}\right\}$ (A.52)
A.3 Key trigonometric identities
Finally we provide the key trigonometric identities we will use in this paper when dealing with Fourier series of trading strategies.
$\int_{0}^{1}\sin(n\pi t)\,\texttt{dt}$ =$=\begin{cases}\frac{2}{n\pi}&\text{if $n$ is odd}\\ 0&\text{otherwise}\end{cases}$ (A.53a)
$\int_{0}^{1}\cos(n\pi t)\,\texttt{dt}$ =$=\begin{cases}1&\text{if }n=0\\ 0&\text{if }n\neq 0\end{cases}$ (A.53b)
$\int_{0}^{1}t\cos(n\pi t)\,\texttt{dt}$ =$=\begin{cases}-\frac{2}{n^{2}\pi^{2}}&\text{if $n$ is odd}\\ 0&\text{otherwise}\end{cases}$ (A.53c)
$\int_{0}^{1}\cos(n\pi t)\cos(m\pi t)\,\texttt{dt}$ =$=\begin{cases}\frac{1}{2}&\text{if }n=m\\ 0&\text{if }n\neq m\\ \end{cases}$ (A.53d)
[…]
Appendix B Acknowledgements
I extend my sincere thanks to Vladimir Ragulin of Manifold Insights for building the quadratic optimization code for solving the constrained and unconstrained optimizations, finding two-trader equilibria and producing the plots in this paper. He also introduced the use of state-space diagrams which have turned out to shed additional light on the nature of equilibria. His careful reading of drafts of this paper and transcription to code identified many extremely valuable insights and improvements in the work.